OpenStudy (anonymous):

Find the product of (x+3)(x+5) and explain the solution..

OpenStudy (mimi_x3):

x = -3 or x=-5

OpenStudy (chaise):

x * x = x^2 x * 5 = 5x 3 * x = 3x 3 x 5 = 15 x^2 + 8x +15 Do you see how I multiplied out the bracket? I hope this is what you asked for, and not the solution to your equation.

OpenStudy (mimi_x3):

lols , woops i misunderstood the question xD

OpenStudy (anonymous):

THANKS CHAISE

OpenStudy (anonymous):